OpenStudy (anonymous):

The sum of the angles of a triangle is 180o. The first and second angles are equal. The third angle is 2.5 times the first. How large is the first angle?

OpenStudy (anonymous):

Like do that say 180 or 1800

OpenStudy (anonymous):

it says 180 degrees with the little circle symbol

OpenStudy (anonymous):

alright lets set up the equation. x+x+2.5x=180 becuase two angles are the same (x) and the third angle is 2.5 times the first (2.5x). if we simplify the equation we get 4.5x=180 next we divide both sides by 4.5, the answer is x=40 so the first two angles are 40, and the largest angle is 40 times 2.5, which is 100. we know this is true because 40+40+100=180

OpenStudy (anonymous):

@DUDE..LOOK..ITS..SGTC any questions?

OpenStudy (anonymous):

Find the revenue from selling 22 items if the revenue function is R(x) = 250 − 7x. 102 211 96 243

OpenStudy (anonymous):

@moudown828

OpenStudy (anonymous):

wouldnt be like 96